Gemini completed these transactions during 2015:

March 1 The directors declared a $0.75 per share cash dividend payable on March 31, to the March 15 shareholders of record.
June 1 Issued 12,000 common shares at $6.50 per share.
October 1 Issued 8,000 common shares at $4.75 per share.
December 1 The directors declared a $1.25 per share cash dividend payable on January 2, 2016, to the December 15 shareholders of record.
December 31 Net income for the year is $245,000.
Required:
2. Prepare the equity section of Gemini’s balance sheet at December 31, 2015.
1
Expert's answer
2020-02-27T09:34:00-0500

Balance at the beginning:

500 000 (100 000 shares)

400 000 net profit

Coming:

12 000*6.50+8 000*4.50=78 000+38 000=116 000 (20 000)

Consumption

0.75*100 000+1.25*100 000=75 000+125 000=200 000

Balance at the end

616 000(120 000 shares)

200 000 net profit